McConnell Corporation has bonds on the market with 20 years to maturity, a YTM of 6.8 percent, a par value of $1,000, and a current price of $1,326.50. The bonds make semiannual payments. What must the coupon rate be on these bonds?
Answer #1

Information provided:

Par value= Future value= $1,000

Present value= $1,326.50

Time= 20 years*2= 40 semi-annual periods

Yield to maturity= 6.80%/2= 3.40% per semi-annual period

The coupon rate is calculated by first computing the coupon payments.

The coupon payment is calculated by entering the below in a financial calculator:

FV= 1,000

PV= -1,326.50

N= 40

I/Y= 3.40

Press the CPT key and PMT to compute the amount of coupon payment.

The value obtained is $49.0528 per semi-annual period.

Therefore, the amount of annual coupon payment is $49.0528*2= $98.11.

Coupon rate is calculated using the below formula:

Coupon rate= Annual coupon payment/ Par value of bond*100

                    = $98.11/ $1,000*100

                    = 0.0981*1,000

                    = 9.81%.

The coupon rate of the bond is 9.81%.